# United Minds or Isolated Agents? Exploring Coordination of LLMs under Cognitive Load Theory

### v1

@@ 4 CoThinker
CoThinker is a multi-agent ICL architecture designed to enhance collaborative problem-solving by systematically managing cognitive load. Simply aggregating outputs from LLM agents often proves insufficient for complex tasks, as naive collaboration can introduce significant transactional costs—the cognitive effort required to coordinate, communicate, and integrate—without a corresponding increase in solution quality (Pan et al., 2025). As Cognitive Load Theory (CLT) suggests, these transactional costs can quickly lead to extraneous cognitive overload, negating the benefits of parallel thinking (Kirschner et al., 2009, 2018). To overcome these challenges within the ICL paradigm, CoThinker operationalizes the principles of human collective intelligence discussed in Section 3, aiming to create a "collective mind" that distributes cognitive load. We leverage the insights from CLT to design an architecture that mirrors how human groups effectively solve complex problems.
Figure 2: The CoThinker Architecture. A high Cognitive Load (CL) task is initially processed by diverse agents via Agent Parallel Thinking. The Transactive Memory System (TMS) facilitates shared understanding by updating and allowing retrieval of collective knowledge. The Communication Moderator manages inter-agent information flow, leveraging a trade-off to form a cognitive small-world network, which then feeds into the Synthesizer to produce a final solution, resulting in a lower effective CL for the overall system.
To operationalize these insights, the CoThinker architecture (Figure 2) comprises four main modules: Agent Parallel Thinking (Section 4.1), Transactive Memory System (Section 4.2), Communication Moderator (Section 4.3), and Synthesizer (Section 4.4). Each module is directly guided by CLT principles to emulate aspects of the human collective mind. Agent Parallel Thinking fosters initial cognitive diversity, potentially splitting the intrinsic load of the task. The Transactive Memory System boosts inter-agent understanding and tracks consensus, reducing cognitive load from redundant processing. The Communication Moderator balances intrinsic and extraneous loads by structuring information exchange. Finally, the Synthesizer integrates refined collective insights. Let $\mathcal{A}=\{A_{1},\dots,A_{M}\}$ be the set of $M$ agents. Let $T_{\text{max}}$ be the total number of generation rounds. Agent $A_{i}$’s output at the end of round $t$ is denoted $x_{i}^{(t)}$.
@@ 4.1 Agent Parallel Thinking
This module promotes a division of cognitive labor and parallel thinking by assigning diverse thinking styles. Unlike assigning pre-defined roles, which require domain-specific foresight and impose extraneous cognitive load from role adherence, CoThinker uses an adaptive approach. A Thinking Style Orchestrator generates a task-specific style $\phi_{i}$ for each agent $A_{i}$ based on a general base thinking style inventory $\psi$ (Sternberg, 1997) and the task $D$:
@@
This yields diverse thinking styles $\{\phi_{i}\}_{i=1}^{M}$, employed in subsequent stages. Further details on the prompting strategy for style generation and thinking style inventory are in the Appendix (B.1).
@@ 4.2 Transactive Memory System (TMS)
Human groups effectively manage complex information by developing Transactive Memory Systems (TMS), which involve a shared understanding of who knows what, how to access information held by others, and a collective agreement on the information itself (Wegner, 1987; Hollingshead, 2001). This distributed cognitive system allows individuals to specialize and rely on others, reducing individual cognitive load and enhancing group problem-solving (Lewis, 2003). To emulate these benefits and foster a collective working memory in CoThinker, we implement a structured mechanism for maintaining and accessing shared knowledge. At each round $t$, an evolving representation of the group’s collective knowledge, denoted $\mu^{(t)}$, is updated based on contributions from all agents:
@@
This aims to enhance shared awareness and efficient integration of distributed knowledge. The specific prompt-based emulation of TMS components is detailed in the Appendix (See B.2).
@@ 4.3 Communication Moderator
Effective inter-agent communication is crucial, yet it incurs transactional costs—the cognitive effort for message processing and integration—which can impose extraneous cognitive load, a key concern in Collaborative Cognitive Load Theory (Kirschner et al., 2009, 2018). To mitigate these costs, the Communication Moderator structures information exchange by selecting $N$ reference messages $\mathcal{P}_{i}^{(t-1)}$ for each agent $A_{i}$. This process navigates the critical trade-offs between Network Density vs. Sparsity (high exposure and cost vs. low cost and potential information loss) and Information Homogeneity vs. Heterogeneity. The latter involves balancing the ease of integrating cognitively similar inputs (low extraneous load but risk of echo chambers (Runkel, 1956)) against the benefits of diverse perspectives for distributing intrinsic load (Aral and Van Alstyne, 2011)).
Communication Topology and Algorithm: The selection of references defines a directed communication graph $G^{(t-1)}=(\mathcal{A},E^{(t-1)})$ for each round, where an edge $(A_{u},A_{v})\in E^{(t-1)}$ exists if agent $A_{v}$ receives a message from agent $A_{u}$ generated in round $t-1$. Motivated by how small-world networks efficiently balance local clustering with global connectivity (Watts and Strogatz, 1998), our moderator employs the following algorithm to construct this graph:
@@
    Set Fixed In-Degree ($N$): Each agent $A_{i}$ (node $A_{v}$) has an in-degree of $N$, capping its processing load and respecting LLM working memory (Zhang et al., 2024b; Gong et al., 2024).
@@
    Connection Establishment via Probabilistic Rewiring ($\beta$): For each agent $A_{i}$, its $N$ incoming edges (references $\mathcal{P}_{i}^{(t-1)}$) are established by primarily choosing messages from cognitively similar peers (low distance), but with a probability $\beta$, "rewiring" some connections to randomly chosen, diverse peers.
Resulting Network Properties and Cognitive Balance: This rewiring process fosters dynamic communication networks $G^{(t-1)}$ with small-world properties. Such networks exhibit high local clustering (facilitating efficient refinement of similar ideas, reducing extraneous load locally) and short average path lengths (enabling rapid global propagation of diverse insights, aiding intrinsic load distribution). This structure offers a principled balance between focused collaboration and broad information access, managing cognitive load more effectively than purely random or regular lattice networks.
@@ 4.4 Synthesizer
The Synthesizer consolidates information into a final solution after $T_{max}$ rounds. It can be an External Agent (dedicated LLM) or an In-group Agent (team member) (Lu et al., 2024; Shinn et al., 2023). This draws from Collaborative Cognitive Load Theory (Kirschner et al., 2018) and Observational Learning (Bandura and Walters, 1977) (See B.4)
CoThinker Process Flow
The process for task $D$ with $M$ agents over $T$ rounds:
@@
Iterative Refinement: For each agent $A_{i}$ and round $t$:
$\mathcal{P}_{i}^{(t)}$ $=\operatorname{SelectRefs}\bigl{(}\{x_{k}^{(t)}\}_{k\in\mathcal{A}\setminus\{A_{i}\}},x_{i}^{(t)},N,\beta\bigr{)}$ (4)
$x_{i}^{(t+1)}$ $=\operatorname{Agent}\bigl{(}D,\phi_{i},\mu^{(t)},x_{i}^{(t)},\mathcal{P}_{i}^{(t)}\bigr{)}$ (5)
$\mu^{(t+1)}$ $=\operatorname{UpdateMem}\bigl{(}\mu^{(t)},\{x_{i}^{(t+1)}\}_{i=1}^{M}\bigr{)}$ (6)
@@ 5 Experiments and Results
This section details our experimental methodology and presents the empirical evaluation of CoThinker. We first outline the experimental setup, including the base LLMs, benchmarks, and baselines. We then present the main results on LiveBench and CommonGen-Hard, followed by ablation studies and a discussion of our findings through the lens of Cognitive Load Theory (CLT).
@@ 5.1 Experimental Setup
Models and Configuration. We use three Gemini models (Team et al., 2024) with varying capacities: gemini-1.5-flash-8b (lightweight), gemini-1.5-flash (mid-tier), and gemini-1.5-pro (high-capacity). All models run with the initial generation temperature set to 0.25 to encourage diverse outputs. In multi-agent settings, subsequent rounds use temperature 0.0 and a frequency penalty of 0.5 to reduce repetition. By default, multi-agent methods use $M{=}6$ agents interacting over $T{=}3$ rounds. For CoThinker, we set $N{=}3$ references and exploration parameter $\beta{=}0.3$.
Evaluation Benchmarks. We evaluate on two challenging benchmarks: (1) LiveBench (White et al., 2025), a recent diverse suite drawing from Big-Bench Hard (Suzgun et al., 2023), AMPS (Hendrycks et al., 2021), and IFEval (Zhou et al., 2023), covering domains such as mathematics, coding, language, instruction following, and data analysis; and (2) CommonGen-Hard (Madaan et al., 2023), a cognitively demanding variant of CommonGen (Lin et al., 2020), which evaluates multi-sentence generation under high element interactivity. We adopt a 10-dimensional metric for CommonGen-Hard evaluation (Li et al., 2018) (See C.1).
Baselines. We compare CoThinker with both single-agent and multi-agent approaches. (i) Single Agent (IO) is a standard mode of prompting. (ii) Single Agent (CoT) incorporates Chain-of-Thought prompting (Wei et al., 2022). (iii) Single Agent (Self-Refine) uses iterative self-critique and revision (Madaan et al., 2023). (iv) Multi-Agent Debate (MAD): employs interactive agent discussion with consensus formation (Du et al., 2023; Liang et al., 2023). (v) Diverse MAD (DMAD): introduces heterogeneous prompting to avoid fixed mental sets (Liu et al., 2025b) (See C.2)
@@ 5.2 Main Results on LiveBench
Table (1) presents the performance of CoThinker and baseline methods across the LiveBench suit for gemini-1.5-flash-8b, gemini-1.5-flash, and gemini-1.5-pro. Scores are reported as relative improvements over the respective gemini-8b-flash’s IO (Standard Prompt) baseline. An average score is calculated as the arithmetic mean of these relative scores across the main LiveBench categories (Math, Reasoning, Instruction, Data, Language).
@@
Method Math Data Reas. Lang. Instr. Avg. Math Data Reas. Lang. Instr. Avg. Math Data Reas. Lang. Instr. Avg.
IO 1.00 1.00 1.00 1.00 1.00 1.00 1.47 2.03 1.63 1.41 1.10 1.53 2.00 2.92 1.87 1.43 1.03 1.85
CoT 1.04 0.90 1.11 1.09 1.02 1.03 1.47 2.07 1.74 1.30 1.10 1.54 1.86 2.72 1.82 1.54 1.02 1.79
SR 0.92 0.34 0.80 0.89 0.81 0.75 1.45 0.90 1.55 1.06 0.87 1.17 1.93 1.33 1.80 1.22 0.72 1.40
MAD 1.13 0.58 1.21 1.03 0.87 0.97 1.51 1.46 1.92 1.46 1.01 1.47 2.29 3.15 1.78 1.58 0.77 1.92
DMAD 1.13 0.64 0.85 1.02 0.89 0.91 1.49 2.51 1.94 1.44 1.06 1.69 2.31 3.32 1.88 1.74 1.02 2.05
CoThinker 1.11 1.32 1.22 0.98 0.80 1.07 1.57 2.44 1.97 1.52 0.99 1.70 2.40 3.39 1.95 1.76 0.95 2.09
Table 1: LiveBench(White et al., 2025) performance with all scores normalized by gemini-1.5-flash-8b-io baseline. The abbreviations corresponded to Math, Data Analysis, Reasoning Language, and Instruction Following
Analysis of LiveBench Results. CoThinker consistently achieves strong average performance across all base model families, with particularly notable gains in complex categories like Data Analysis, Reasoning, and often Math, but low performance on Instruction Following. We posit this performance pattern reflects two distinct task categories: those with high intrinsic cognitive load and those with low intrinsic load. The former, characterized by tasks like Data Analysis and Reasoning, demonstrates a clear scaling in baseline performance as model capability increases (e.g., from gemini-1.5-flash-8b to gemini-1.5-pro), indicating that greater raw cognitive power inherently improves outcomes. For these high-load tasks, CoThinker excels by effectively decomposing complex problems and orchestrating collaborative agent contributions, therefore, splitting the intrinsic load to enhance performance.
Conversely, tasks with low intrinsic load, such as instruction following (Instr.), show minimal or inconsistent performance gains when moving from weaker to stronger base models; for instance, the gemini-1.5-pro IO baseline on Instruction Following does not substantially outperform that of gemini-1.5-flash-8b. This suggests the primary bottleneck is not cognitive load. In such scenarios, the added communication overhead inherent in CoThinker can outweigh the benefits of collaboration. For tasks demanding straightforward adherence rather than sophisticated reasoning, this introduced more extraneous cognitive load, explaining why CoThinker may not show an advantage or can even underperform on these low-load, execution-focused tasks.
@@ 5.3 Main Results on CommonGen-Hard
In CommonGen-Hard, which emphasizes managing high element interactivity, CoThinker demonstrates notable performance improvements. Figure 3 presents these results, with Figure 3(a) illustrating its balanced strengths across evaluation dimensions and Figure 3(b) showing performance trends over interaction rounds.
@@
Figure 3: CoThinker performance on CommonGen-Hard using gemini-1.5-flash. (a) The radar plot illustrates a multi-dimensional performance profile, where CoThinker typically shows well-rounded and superior strengths compared to baselines. (b) The rounds plot depicts the total score trend across interaction rounds ($T$), often indicating an optimal number of rounds before performance plateaus or declines.
Analysis of CommonGen-Hard Results.
CoThinker demonstrates strong overall performance on CommonGen-Hard (Figure 3), effectively managing its high element interactivity. The multi-dimensional profile (Figure 3(a)) typically shows CoThinker excelling in key areas like coherence and concept integration, albeit with occasional trade-offs in aspects such as conciseness. This aligns with Cognitive Load Theory (CLT); the high intrinsic load of the task is managed by CoThinker’s distributed processing and transactive memory. Notably, its performance trajectory versus interaction rounds (Figure 3(b)) highlights a key advantage: CoThinker achieves sustained constructive refinement over several rounds, effectively managing cognitive load. This contrasts with the multi-agent baseline where performance degrades due to rapidly accumulating extraneous load from inefficient coordination or information overload. CoThinker’s architecture appears more adept at balancing these loads, delaying diminishing returns.
5.4 Ablation Studies on LiveBench Subsets
Ablation studies were conducted on gemini-1.5-flash-8b using averaged scores from selected LiveBench subtask categories (Math, Reasoning, Data Analysis, and Instruction). These studies investigated the impact of CoThinker’s reference set size ($N$), exploration rate ($\beta$), and the number of agents ($M$). Unless otherwise specified, default parameters were $T=3$. For $N$ ablation, $M=6,\beta=0.3$. For $\beta$ ablation, $N=2,M=6$. For $M$ ablation, $N=3,\beta=0.3$. All scores are normalized by the I/O baseline performance for each subtask before category averaging.
Figure 4: Ablation studies on CoThinker parameters ($N,\beta,M$) using gemini-1.5-flash-8b. Performance is shown for four LiveBench task categories (Math, Reasoning, Data Analysis, Instruction), normalized by IO baseline performance (1.0). Top row: Effect of Reference Set Size ($N$), varying $N\in\{0,2,3,4,5\}$ with $M=6,\beta=0.3,T=3$. Middle row: Effect of Exploration Rate ($\beta$), varying $\beta\in\{0.1,0.3,0.6,1.0\}$ with $N=2,M=6,T=3$. Bottom row: Effect of Number of Agents ($M$), varying $M\in\{6,12,18\}$ with $N=3,\beta=0.3,T=3$. Optimal parameter settings are task-dependent, indicating varying sensitivities to peer input diversity and information overload.
Analysis of Ablation Studies.

### v2

@@ 4 CoThinker
CoThinker is the operationalization of the CLT principles and collective intelligence mechanisms outlined in Section 3, designed as a multi-agent ICL architecture that systematically manages CL to enhance collaborative problem-solving. Simply aggregating outputs from LLM agents often proves insufficient for complex tasks, as naive collaboration can introduce transactional costs—the cognitive effort required to coordinate, communicate, and integrate—without a corresponding increase in solution quality (Pan et al., 2025). As CLT predicts, these transactional costs can quickly lead to cognitive overload, negating the benefits of parallel thinking (Kirschner et al., 2009; 2018). CoThinker addresses these challenges by translating the three core CLT principles into a practical multi-agent framework that creates a "collective mind" capable of distributing CL effectively.
Figure 2: The CoThinker Architecture. A high CL task is initially processed by diverse agents via Agent Parallel Thinking. Transactive Memory System facilitates shared understanding by updating and retrieving of collective knowledge. Communication Moderator manages inter-agent information flow, leveraging a trade-off to form a cognitive small-world network, which then feeds into the Synthesizer for final solution, resulting in a lower effective CL for the overall system.
To operationalize these insights, the CoThinker architecture (Figure 2) comprises four main modules: Agent Parallel Thinking (Section 4.1), Transactive Memory System (Section 4.2), Communication Moderator (Section 4.3), and Synthesizer (Section 4.4). Each module is directly guided by CLT principles to emulate aspects of the human collective mind. Agent Parallel Thinking fosters initial cognitive diversity, potentially splitting the intrinsic load of the task. The Transactive Memory System boosts inter-agent understanding and tracks consensus, reducing CL from redundant processing. The Communication Moderator balances intrinsic and extraneous loads by structuring information exchange. Finally, the Synthesizer integrates refined collective insights. Let $\mathcal{A}=\{A_{1},\dots,A_{M}\}$ be the set of $M$ agents. Let $T_{\text{max}}$ be the total number of generation rounds. Agent $A_{i}$’s output at the end of round $t$ is denoted $x_{i}^{(t)}$.
@@ 4.1 Agent Parallel Thinking
This module promotes a division of cognitive labor and parallel thinking by assigning diverse thinking styles. Unlike assigning pre-defined roles, which require domain-specific foresight and impose extraneous CL from role adherence, CoThinker uses an adaptive approach. A Thinking Style Orchestrator generates a task-specific style $\phi_{i}$ for each agent $A_{i}$ based on a general base thinking style inventory $\psi$ (Sternberg, 1997) and the task $D$:
@@
This yields diverse thinking styles $\{\phi_{i}\}_{i=1}^{M}$, employed in subsequent stages. Further details on the prompting strategy for style generation and thinking style inventory are in Appendix D.3.
@@ 4.2 Transactive Memory System (TMS)
Human groups effectively manage complex information by developing Transactive Memory Systems (TMS), which involve a shared understanding of who knows what, how to access information held by others, and a collective agreement on the information itself (Wegner, 1987; Hollingshead, 2001). This distributed cognitive system allows individuals to specialize and rely on others, reducing individual CL and enhancing group problem-solving (Lewis, 2003). To emulate these benefits and foster a collective working memory in CoThinker, we implement a structured mechanism for maintaining and accessing shared knowledge. This enables cognitive offloading and specialization. Implementation details in Appendix D.4. At each round $t$, an evolving representation of the group’s collective knowledge, denoted $\mu^{(t)}$, is updated based on contributions from all agents:
@@ 4.3 Communication Moderator
Effective inter-agent communication is crucial, yet it incurs transactional costs, the cognitive effort for message processing and integration, imposing extra extraneous CL. To mitigate these costs, Communication Moderator selects $N<M$ reference messages for each agent $A_{i}$. This process navigates the critical trade-offs between Network Density vs. Sparsity (high exposure vs. information loss) and Information Homogeneity vs. Heterogeneity. The latter involves balancing the ease of integrating cognitively similar inputs (low extraneous load but risk of echo chambers (Runkel, 1956)) against the benefits of diverse perspectives for distributing intrinsic load (Aral & Van Alstyne, 2011).
Communication Topology and Algorithm: The selection of references defines a directed communication graph $G^{(t-1)}=(\mathcal{A},E^{(t-1)})$ for each round, where an edge $(A_{u},A_{v})\in E^{(t-1)}$ exists if agent $A_{v}$ receives a message from agent $A_{u}$ generated in round t1t-1. Motivated by how small-world networks efficiently balance local clustering with global connectivity (Watts & Strogatz, 1998), our moderator employs the following algorithm to construct this graph:
@@
    Fixed In-Degree (NN): Each agent $A_{i}$ (node $A_{v}$) has an in-degree of $N$, capping its processing load and respecting LLM WM (Zhang et al., 2024b; Gong et al., 2024).
@@
    Re-connection via Probabilistic Rewiring ($\beta$): For each agent $A_{i}$, its $N$ incoming edges (references $\mathcal{P}_{i}^{(t-1)}$) are chosen from cognitively similar peers (low distance), but with a probability $\beta$, "rewiring" some connections to randomly chosen, diverse peers.
Resulting Network Properties and Cognitive Balance: It fosters dynamic communication networks with small-world properties, with high local clustering (facilitating efficient refinement of similar ideas, reducing extraneous load locally) and short average path lengths (enabling rapid global propagation of diverse insights, aiding intrinsic load distribution). This structure offers a balance between focused collaboration and broad information access, managing CL more effectively than random or regular lattice networks. Further details are in Appendix D.5.
@@ 4.4 Synthesizer
The Synthesizer consolidate all agents answer and TMS into a final answer (details in Appendix D.6). CoThinker Process Flow. The process for task $D$ with $M$ agents over $T$ rounds:
@@
Iterative Refinement for agent $A_{i}$ and round $t$:
$\mathcal{P}_{i}^{(t)}=\operatorname{SelectRefs}$ $\bigl(\{x_{k}^{(t)}\},N,\beta\bigr)$ (4)
$x_{i}^{(t+1)}=\operatorname{Agent}\bigl(D,\phi_{i},\mu^{(t)},x_{i}^{(t)},\mathcal{P}_{i}^{(t)}\bigr),\quad$ $\mu^{(t+1)}=\operatorname{UpdateMem}\bigl(\mu^{(t)},\{x_{k}^{(t+1)}\}\bigr)$ (5)
@@ 5 Experiments and Results
This section details our experimental methodology and presents the empirical evaluation of CoThinker. We first outline the experimental setup, and then present the main results on LiveBench (White et al., 2025) and CommonGen-Hard (Madaan et al., 2023), followed by ablation studies and a discussion of our findings through the lens of Cognitive Load Theory (CLT).
@@ 5.1 Experimental Setup
Models and Configuration. For main experiments, we use three Gemini models (Team et al., 2024) with varying capacities: Gemini-1.5-Flash-8B (lightweight), Gemini-1.5-Flash (mid-tier), and Gemini-1.5-Pro (high-capacity). Evaluation Benchmarks. We evaluate on two challenging benchmarks: (1) LiveBench (White et al., 2025), a comprehensive SOTA benchmark containing real-world tasks, providing broad-spectrum evaluation (math, reasoning, data analysis and so on); and (2) CommonGen-Hard (Madaan et al., 2023), a controlled experimental challenge designed to test information interactivity, by forcing models to integrate target concepts from large pools of distractors. Baselines. We compare CoThinker with both single-agent and multi-agent approaches: Single Agent (IO), Single Agent (CoT) (Wei et al., 2022), Single Agent (Self-Refine) (Madaan et al., 2023), Multi-Agent Debate (MAD) (Du et al., 2023; Liang et al., 2023), and Diverse MAD (DMAD) (Liu et al., 2025b). Complete details are in Appendices E.2, and E.1.
@@ 5.2 Main Results on LiveBench
Table 2 shows CoThinker performance across three Gemini models. Scores are normalized by the 8B model’s IO baseline. CoThinker achieves strong average performance, excelling in complex tasks (Data Analysis, Reasoning, Math) but underperforming on Instruction Following. This pattern reflects distinct tasks with different cognitive load (CL) characteristics: (1) High intrinsic CL tasks show clear performance scaling with model capability increases, indicating cognitive bottlenecks, whereCoThinker excels by distributing intrinsic CL across agents. (2) Low intrinsic CL tasks show minimal gains from stronger models. CoThinker’s communication overhead introduces extraneous CL that outweighs collaboration benefits, explaining underperformance on execution-focused tasks requiring straightforward adherence rather than reasoning.
@@
Task
IO
CoT
SR
MAD
DMAD
Ours
IO
CoT
SR
MAD
DMAD
Ours
IO
CoT
SR
MAD
DMAD
Ours
Math 1.00 1.04 0.92 1.13 1.13 1.11 1.47 1.47 1.45 1.51 1.49 1.57 2.00 1.86 1.93 2.29 2.31 2.40
Data 1.00 0.90 0.34 0.58 0.64 1.32 2.03 2.07 0.90 1.46 2.51 2.44 2.92 2.72 1.33 3.15 3.32 3.39
Reas. 1.00 1.11 0.80 1.21 0.85 1.22 1.63 1.74 1.55 1.92 1.94 1.97 1.87 1.82 1.80 1.78 1.88 1.95
Lang. 1.00 1.09 0.89 1.03 1.02 0.98 1.41 1.30 1.06 1.46 1.44 1.52 1.43 1.54 1.22 1.58 1.74 1.76
Instr. 1.00 1.02 0.81 0.87 0.89 0.80 1.10 1.10 0.87 1.01 1.06 0.99 1.03 1.02 0.72 0.77 1.02 0.95
Avg. 1.00 1.03 0.75 0.97 0.91 1.07 1.53 1.54 1.17 1.47 1.69 1.70 1.85 1.79 1.40 1.92 2.05 2.09
Table 2: LiveBench (White et al., 2025) performance normalized by Gemini-8B (IO). Ours refers to CoThinker. Abbreviations: Math, Data Analysis, Reasoning, Language, and Instruction Following.
@@ 5.3 Main Results on CommonGen-Hard
CoThinker demonstrates improvements on CommonGen-Hard, which tests high CL management. Figure 3 shows performance across evaluation dimensions through (a) a radar plot with normalized by dividing the min scores and (b) an interaction rounds plot tracking performance evolution. CoThinker effectively handles high element interactivity by distributing CL across specialized agents and leveraging transactive memory—core CLT principles in action. The radar plot (Figure 3(a)) reveals CoThinker’s strengths in coherence and concept integration, with minor trade-offs in conciseness. The rounds plot (Figure 3(b)) shows sustained improvement across multiple interaction rounds: while baseline methods degrade due to accumulated coordination overhead (extraneous CL).
@@
Figure 3: CoThinker performance on CommonGen-Hard (Madaan et al., 2023) using Gemini-1.5-Flash. (a) The radar plot illustrates a multi-dimensional performance, showing well-rounded improvement. (b) The rounds plot depicts the total score across rounds, showing stable improvement.
5.4 Cross-Model Generalization
To validate CoThinker’s model-agnostic nature, we evaluated across multiple LLM families: GPT-5-Nano, Qwen3-30B-A3B, GPT-OSS-20B, Gemini-2.5-Flash, GPT-4.1-Mini, Qwen3-32B, and DeepSeek-R1-8B (partial results in Table 3). We examine two scenarios: (1) standard setting with IO baselines using maximum reasoning steps with temperature 0.25 and (2) constrained setting with token budget of 8192 with greedy decoding (temperature = 0). Complete results are in Appendix E.3.
Standard Setting Model Method Math Reason. GPT-5 CoThinker 88.57 81.88 IO 82.63 68.38 Qwen3 CoThinker 80.62 89.50 IO 77.50 76.00
Constrained Setting Model Method Math Reason. Gemini-2.5 CoThinker 76.3 69.2 IO 59.3 31.0 GPT-4.1 CoThinker 40.0 70.8 IO 34.0 40.8
Table 3: Cross-model evaluation on LiveBench (White et al., 2025) Math and Reasoning (Reason.) subsets in standard setting (left) and constrained setting (right).
5.5 CoThinker Ablation of Communication Moderator